# HLE-Verified: A Systematic Verification and Structured Revision of Humanity’s Last Exam

### v2

@@ Abstract
Humanity’s Last Exam (HLE) has become a widely used benchmark for evaluating frontier large language models on challenging, multi-domain questions. However, subsequent community-led analyses have raised concerns that HLE contains a non-trivial number of noisy items (e.g., ambiguous statements, incorrect answers, or mismatched rationales), which can systematically bias evaluation results and distort cross-model comparisons. To address this challenge, we introduce HLE-Verified, a verified and revised version of HLE accompanied by a transparent, component-wise verification protocol and fine-grained error taxonomy. Our construction follows a two-stage validation-and-repair workflow resulting in a unified certified benchmark. In Stage I, each item is subjected to binary validation on the problem and final answer dimensions (with rationale used as an auxiliary consistency signal), combining domain-expert review and model-based cross-checks. This stage yields 641 items verified as correct. In Stage II, items identified as flawed but fixable are systematically revised under strict constraints that preserve the original evaluation intent, through dual independent expert repairs, model-assisted consistency auditing, and final expert adjudication, resulting in 1,170 revised-and-certified items. The remaining 689 items are released as a documented uncertain set with explicit uncertainty sources and required expertise tags for future community refinement. We compare the performance of seven state-of-the-art language models on HLE and HLE-Verified, observing an average absolute accuracy gain of 7–10 percentage points on HLE-Verified. The improvement is particularly pronounced on items where the original HLE problem statement and/or reference answer is erroneous: on this subset, the models achieve an average accuracy increase of 30–40 percentage points. Moreover, our analyses indicate a strong association between model confidence and the presence of errors in the problem statement or reference answer, providing evidence for the effectiveness of our revisions. Overall, HLE-Verified improves HLE-style evaluations by reducing annotation noise and enabling more faithful measurements of model capabilities. Data is available at: https://github.com/SKYLENAGE-AI/HLE-Verified.
@@ 1 Introduction
As frontier language models (LMs) continue to advance rapidly, there is increasing demand for evaluation benchmarks that are simultaneously difficult, broad in disciplinary coverage, and resistant to saturation. High-difficulty benchmarks play a central role in tracking progress, informing capability claims, and shaping scientific discourse about model reasoning abilities. In this context, Humanity’s Last Exam (HLE) has emerged as a widely used benchmark for assessing model performance on challenging, multi-domain questions spanning mathematics, science, engineering, and the humanities (Center for AI Safety et al., 2026). Results on HLE have been cited to support claims regarding reasoning depth, generalization capacity, and reliability of state-of-the-art models.
However, as benchmark difficulty increases, so does the importance of annotation integrity. When many items lie near a model’s decision boundary, even minor inconsistencies in problem statements or reference answers can disproportionately influence aggregate metrics. Recent independent audits and community-led examinations have raised concerns that HLE contains a non-trivial number of noisy items, including ambiguous or underspecified problem statements, incorrect reference answers, and inconsistencies between rationales and final answers (A. White (2025); lhl (2026); 15). These analyses suggest that a portion of measured model performance on HLE may reflect annotation artifacts rather than genuine capability differences.
@@
To address this reliability challenge, we construct HLE-Verified, a rigorously verified and partially revised version of HLE developed under a structured two-stage verification and revision protocol. Our approach adopts a component-wise validation framework that treats the problem statement and final answer as primary objects of correctness assessment, while using the rationale as a secondary signal for detecting internal inconsistency or under-specification. Under this protocol, HLE-Verified is organized into three disjoint subsets: verified gold items, revised items corrected under preserved evaluation objectives, and uncertain items whose validity cannot be conclusively determined.
We position HLE-Verified as dataset infrastructure: not a new benchmark task, but a methodological reliability enhancement that enables more rigorous, interpretable, and reproducible model comparisons. In addition, we provide an empirical evaluation of how verification and revision affect downstream model measurements, and we release structured metadata that supports continued community-driven refinement.
@@
  • We release HLE-Verified, a verified and revised version of HLE constructed via a transparent component-wise verification protocol and fine-grained error taxonomy, comprising 641 verified items, 1,170 revised-and-verified items, and a 689-item documented uncertain set for structured community refinement.
@@
  • We benchmark seven state-of-the-art LLMs on HLE vs. HLE-Verified, demonstrating that verification materially alters measured performance (an average +7–10 accuracy points overall and +30–40 points on items with erroneous problems/answers), and we analyze the relationship between model confidence and problem/answer errors.
@@ 3.1 Overview of the pipeline
    Gold subset (641 items): validated without modification.
@@
    Revision subset (1,170 items): corrected and re-verified under preserved evaluation objectives.
@@ 3.2 Stage I: component-wise verification
  • Rationale validity (diagnostic): the reference solution is mathematically/logically sound, internally consistent, and compatible with the final answer (when the rationale is intended as an authoritative reference).
@@
Stage I yields the gold (verified) subset of 641 items that are retained without modification. All remaining items are carried forward either to Stage II for repair (if flawed but repairable) or to the uncertain pool (if indeterminate or high-risk).
@@ 3.3 Stage II: systematic revision and re-verification
To ensure that revision results are objectively reviewable, Stage II focuses on domains where correctness can be reliably adjudicated and independently reproduced (e.g., mathematics, physics, chemistry, biomedicine, and computer science in our current release). Items in domains with weak verification boundaries or subjective conventions are handled conservatively (often routed to the uncertain subset unless resolvable with high confidence).
@@
Stage II yields the revision subset of 1,170 items that are corrected and re-verified as suitable for evaluation. Each revised item includes structured revision metadata, including which components were changed, what error types were addressed, and brief adjudication notes sufficient for auditing.
@@ 4.1 Component-wise annotation outcomes
Across all components, valid remains the majority label, indicating that many problematic instances are only partially flawed rather than globally invalid. However, reliability varies systematically by component. The Problem component exhibits the smallest invalid proportion, suggesting that task statements are often structurally well-formed even when other parts of the item are defective. In contrast, the Answer component shows a substantially higher invalid share, reflecting frequent mismatches between the provided answer and the intended solution. The Rationale component is the most error-prone, with the largest invalid proportion overall.
@@ 4.2 Cross-domain variation
For the Problem component, Math and Biology/Medicine exhibit very high validity rates (above 90%), while Chemistry remains comparatively strong (approximately 75% valid). In contrast, Physics, Humanities/Social Science, Other, and Engineering display substantially lower problem validity (approximately 30–35%), with the residual mass dominated by uncertain rather than invalid. This pattern suggests that many questions in these domains are not necessarily incorrect but are harder to verify unambiguously due to domain conventions, interpretive flexibility, or reliance on external knowledge.
For the Answer component, domain gaps persist. Chemistry and Computer Science/AI maintain relatively high answer validity (around 65–70%), whereas Physics, Humanities/Social Science, Other, and Engineering again cluster near 30–35% valid with substantial uncertainty. Notably, Math and Biology/Medicine exhibit a different pattern: uncertainty is comparatively low, but the invalid share is high (roughly 35–40% invalid). This indicates that answer correctness in these domains is often decidable and frequently incorrect rather than ambiguous.
The divergence is most pronounced for the Rationale component. In Math and especially Biology/Medicine, rationales are predominantly labeled invalid, despite the high validity of the corresponding problems. By contrast, in Physics, Humanities/Social Science, Other, and Engineering, rationale defects are more frequently categorized as uncertain, reflecting verification difficulty rather than explicit logical contradiction.
@@ 4.3 Component-wise Defect Distribution Across Subjects
Answer-level defects. Among all answer-level annotations, Incorrect Answer overwhelmingly dominates (approximately 90% of answer defects), while Incomplete, Ambiguous, and format-level answer errors together account for only a small fraction. This indicates that answer failures are predominantly deterministic correctness violations rather than ambiguity or partial specification.
Rationale-level defects. Rationale defects concentrate primarily in structurally incomplete reasoning, especially Missing Prerequisite and related incompleteness categories, along with Format Semantic Error. Classical logical fallacies such as circular reasoning or domain misapplication constitute a comparatively small share. This suggests that the principal weakness of reference rationales lies in structural under-specification rather than overt logical contradiction.
Problem-level defects. Problem-level defects are relatively sparse and are mainly distributed between Format Semantic Error and Missing Information. The near absence of Theoretical Invalidity indicates that most flawed items do not originate from fundamentally incorrect task concepts, but from specification-level imprecision.
Taken together, these global proportions suggest that the primary reliability bottlenecks in HLE arise from answer-key correctness and incomplete rationale construction, rather than from conceptual invalidity of problem statements.
@@
    Deterministic answer errors dominate answer-level defects. Incorrect final answers constitute the principal failure mode across all quantitative disciplines.
@@
    Rationale incompleteness outweighs logical fallacies. Missing assumptions and omitted intermediate steps are substantially more prevalent than classical reasoning contradictions.
@@
    Format-induced semantic distortion is domain-sensitive. Symbolic and notational misalignment contributes meaningfully in Chemistry and Computer Science.
@@
    Conceptual invalidity is rare. Theoretical invalidity forms only a negligible fraction of total defects.
These findings demonstrate that HLE’s problematicness is structured rather than random. Reliability risks are component-specific and discipline-sensitive, with answer correctness and rationale completeness constituting the dominant bottlenecks. This empirical structure justifies a component-wise verification protocol and cautions against treating benchmark noise as uniform or negligible.
@@
Physics Information Missing (57.1%) Information Missing (42.9%) Incorrect Answer (80.0%)
Chemistry Format Semantic Error (58.3%) Format Semantic Error (40.5%) Incorrect Answer (69.4%)
Biology/Medicine Knowledge Error (33.3%) Information Missing (41.9%) Incorrect Answer (97.2%)
Computer Science Format Semantic Error (92.3%) Format Semantic Error (63.4%) Incorrect Answer (78.1%)
@@ 5.1 Experimental Setup
We evaluate the following frontier models: GPT-5.2-Thinking (OpenAI, 2025), Gemini3-Pro-Preview (Google DeepMind, 2025), Claude-Opus4.5 (Anthropic, 2025), Claude-Opus4.6 (Anthropic, 2026), Grok-4.1 (fast-reasoning) (xAI, 2025), DeepSeek-V3.2-Thinking (Liu et al., 2025), and Qwen3-Max-Thinking (Team, 2026). All models were evaluated using the same system prompt recommended by the HLE official guidelines, along with each model’s own default recommended decoding configuration. To reduce variance from stochastic decoding, we run five independent rollouts per item and report avg5 accuracy, i.e., the average correctness across the five sampled completions (without additional post-hoc selection unless otherwise stated).
@@
We report results on (i) the Full Set (Raw HLE Full vs. Revised HLE-Verified Full) to quantify the end-to-end impact on the benchmark, and (ii) a Revised Subset comparison (Raw HLE Subset vs. Revised HLE-Verified Subset) limited to items that were edited/flagged during our verification process, which isolates the effect of flawed items and their corrections. Note that, to ensure comprehensive revisions, we also corrected the original rationales and released them to the open-source community. However, under the standard HLE benchmark evaluation, items with errors only in the rationale do not affect the final scores, since mainstream evaluation uses only the problem and answer. Therefore, the Revised Subset in this experimental section includes only items for which at least one of the problem/answer fields was revised, totaling 624 items (including 572 text-only items). Unless otherwise specified, all evaluation sets used in this section are the text-only subset of HLE.
@@
Claude-Opus4.5 30 55 38.8 46 8.8 14.2 70 48.16 39 33.96
Grok 4.1-fast-reasoning 19.94 73 29 63 9.06 8.25 83 43.07 47 34.82
@@
Deepseek-V3.2 24.9 56 36.4 46 11.5 7.87 70 47.45 28 39.58
Qwen3-Max-Thinking 30 66 38.2 57 8.2 14.2 79 48.48 44 34.28
@@ A.1 LLM Judge Prompt
A.2 LLM Judge Prompt in Stage I
@@
A.3 LLM Judge Prompt in Stage II

### v3

# arXiv:2602.13964v3[cs.CL] 27 Feb 2026

@@ Abstract
Humanity’s Last Exam (HLE) has become a widely used benchmark for evaluating frontier large language models on challenging, multi-domain questions. However, subsequent community-led analyses have raised concerns that HLE contains a non-trivial number of noisy items (e.g., ambiguous statements, incorrect answers, or mismatched rationales), which can systematically bias evaluation results and distort cross-model comparisons. To address this challenge, we introduce HLE-Verified, a verified and revised version of HLE accompanied by a transparent, component-wise verification protocol and fine-grained error taxonomy. Our construction follows a two-stage validation-and-repair workflow resulting in a unified certified benchmark. In Stage I, each item is subjected to binary validation on the problem and final answer dimensions (with rationale used as an auxiliary consistency signal), combining domain-expert review and model-based cross-checks. This stage yields 668 items verified as correct. In Stage II, items identified as flawed but fixable are systematically revised under strict constraints that preserve the original evaluation intent, through dual independent expert repairs, model-assisted consistency auditing, and final expert adjudication, resulting in 1,143 revised-and-certified items. The remaining 689 items are released as a documented uncertain set with explicit uncertainty sources and required expertise tags for future community refinement. We compare the performance of eight state-of-the-art language models on HLE and HLE-Verified, observing an average absolute accuracy gain of 7–10 percentage points on HLE-Verified. The improvement is particularly pronounced on items where the original HLE problem statement and/or reference answer is erroneous: on this subset, the models achieve an average accuracy increase of 30–40 percentage points. Moreover, our analyses indicate a strong association between model confidence and the presence of errors in the problem statement or reference answer, providing evidence for the effectiveness of our revisions. Overall, HLE-Verified improves HLE-style evaluations by reducing annotation noise and enabling more faithful measurements of model capabilities. Data is available at: https://huggingface.co/datasets/skylenage/HLE-Verified.
@@ 1 Introduction
As frontier language models (LMs) continue to advance rapidly, there is increasing demand for evaluation benchmarks that are simultaneously difficult, broad in disciplinary coverage, and resistant to saturation. High-difficulty benchmarks play a central role in tracking progress, validating capability claims, and shaping scientific discourse about model reasoning abilities. Against this backdrop, the Humanity’s Last Exam (HLE) has emerged as a widely adopted benchmark for evaluating model performance on challenging questions across mathematics, science, engineering, and the humanities(Center for AI Safety et al., 2026). Results on HLE have been cited to support claims regarding reasoning depth, generalization capacity, and reliability of state-of-the-art models.
However, as benchmark difficulty increases, so does the importance of annotation integrity. When many items lie near a model’s decision boundary, even minor inconsistencies in problem statements or reference answers can disproportionately influence aggregate metrics. Recent independent audits and community-led reviews have highlighted substantial noise in the HLE dataset, including ambiguous or poorly defined question statements, erroneous reference answers, and inconsistencies between rationales and final answers (A. White (2025); lhl (2026); 17). These analyses suggest that a portion of measured model performance on HLE may reflect annotation artifacts rather than genuine capability differences.
@@
To address this reliability challenge, we developed HLE-Verified—a rigorously validated and partially revised version of HLE built under a structured two-stage validation and revision protocol. This approach employs a component-level verification framework, treating problem statements and final answers as primary targets for correctness assessment while utilizing rationale as secondary signals to detect internal contradictions or underdefined specifications. Under this protocol, HLE-Verified is organized into three disjoint subsets: verified gold items, revised items corrected under preserved evaluation objectives, and uncertain items whose validity cannot be conclusively determined.
We position HLE-Verified as dataset infrastructure: not a new benchmark task, but a methodological reliability enhancement that enables more rigorous, interpretable, and reproducible model comparisons. Furthermore, we provide an empirical evaluation of how verification and revision affect downstream model measurements, and we release structured metadata that supports continued community-driven refinement.
@@
  • We release HLE-Verified, a verified and revised version of HLE constructed via a transparent component-wise verification protocol and fine-grained error taxonomy, comprising 668 verified items, 1,143 revised-and-verified items, and a 689-item documented uncertain set for structured community refinement.
@@
  • We benchmark eight state-of-the-art LLMs on HLE vs. HLE-Verified, demonstrating that verification materially alters measured performance (an average +7–10 accuracy points overall and +30–40 points on items with erroneous problems/answers), and we analyze the relationship between model confidence and problem/answer errors.
@@ 3.1 Overview of the pipeline
    Gold subset (668 items): validated without modification.
@@
    Revision subset (1,143 items): corrected and re-verified under preserved evaluation objectives.
@@ 3.2 Stage I: component-wise verification
  • Rationale validity: the reference solution is mathematically/logically sound, internally consistent, and compatible with the final answer (when the rationale is intended as an authoritative reference).
@@
Stage I yields the gold (verified) subset of 668 items that are retained without modification. All remaining items are carried forward either to Stage II for repair (if flawed but repairable) or to the uncertain pool (if indeterminate or high-risk).
@@ 3.3 Stage II: systematic revision and re-verification
To ensure that revision results are objectively reviewable, Stage II focuses on domains where correctness can be reliably adjudicated and independently reproduced (mathematics, physics, chemistry, biomedicine, and computer science in our current release). Items in domains with weak verification boundaries or subjective conventions are handled conservatively (routed to the uncertain subset).
@@
Stage II yields the revision subset of 1,143 items that are corrected and re-verified as suitable for evaluation. Each revised item includes structured revision metadata, including which components were changed, what error types were addressed, and brief adjudication notes sufficient for auditing.
@@ 4.1 Component-wise annotation outcomes
Across components, the Problem field exhibits the highest reliability. The majority of problem statements are structurally valid, and explicit structural errors constitute only a small fraction of cases. This indicates that most instances are well-posed at the task-description level. In contrast, reliability decreases at the Answer level. Only about half of the answers are fully valid, with a substantial portion either incorrect or difficult to verify. This suggests that correctness mismatches and solution inconsistencies are a major source of dataset noise. The most pronounced degradation appears in the Rationale component. Invalid rationales outnumber valid ones, indicating that reasoning chains frequently contain logical gaps, unsupported inferences, or incorrect derivations—even when the corresponding problem statement appears acceptable.
@@ 4.2 Cross-domain variation
For the Problem component, formal scientific domains exhibit markedly higher structural reliability. Math and Biology/Medicine both exceed 92% validity, while Chemistry remains comparatively strong at 73.3%. In contrast, Physics shows only 30.0% valid problem statements, and Engineering, Humanities/Social Science, and Other fall further to the 22–27% range. Notably, in these lower-validity domains, the majority of remaining cases are labeled uncertain (roughly 67–77%), rather than explicitly invalid, indicating verification ambiguity rather than structural error.
At the Answer level, domain gaps persist. Computer Science/AI (65.1%) and Chemistry (67.3%) achieve the highest answer validity rates. Math (59.6%) and Biology/Medicine (58.6%) remain above 50%, but both exhibit substantial invalid proportions (36.8% and 38.2%, respectively), suggesting that answer correctness in these domains is typically decidable and frequently incorrect rather than ambiguous. By contrast, Physics, Engineering, Humanities/Social Science, and Other remain below 30% valid, with uncertainty again dominating (approximately two-thirds to three-quarters of cases).
The divergence becomes most pronounced for the Rationale component. In Math and Biology/Medicine, invalid rationales dominate (65.9% and 66.4%, respectively), and Chemistry reaches an even higher invalid rate of 70.3%. These domains therefore display explicit logical or derivational errors even when problems are well-formed. In contrast, Engineering, Humanities/Social Science, and Other show extremely high uncertainty rates (approximately 73–77%) with virtually no explicit invalidation, reflecting verification difficulty rather than clear logical contradiction. Physics occupies an intermediate position, with high uncertainty (67.0%) but a non-trivial invalid share (10.9%).
@@ 4.3 Component-wise Defect Distribution Across Subjects
Answer-level defects. Across all five domains, Incorrect Answer (type 1) is consistently the dominant answer-level defect. Its within-component proportion ranges from 69.4% (Chemistry) to 97.2% (Biology/Medicine), with all subjects exceeding 70%. This indicates that answer-level failures are primarily deterministic correctness violations rather than ambiguity, partial specification, or formatting issues.
Rationale-level defects. Rationale defects are more heterogeneous across subjects. In Mathematics (40.1%) and Biology/Medicine (45.0%), type 3 defects (structural incompleteness) dominate, whereas Chemistry (40.0%) and Computer Science (63.3%) are led by type 10 defects (format-induced semantic errors). Physics shows a flatter distribution, with type 3 accounting for 25.0% of rationale defects. Overall, rationale instability arises primarily from missing intermediate structure and representation misalignment rather than classical logical contradictions.
Problem-level defects. Problem-level defect patterns are domain-sensitive. Mathematics (40.0%), Chemistry (56.5%), and Computer Science (94.6%) are dominated by type 5 defects (format semantic errors), whereas Physics is led by type 1 (57.1%) and Biology/Medicine by type 2 (33.3%). The near absence of dominant theoretical-invalidity categories suggests that most flawed items originate from specification or representation-level imprecision rather than fundamentally incorrect task concepts.
Taken together, these proportions indicate that reliability degradation in HLE is primarily driven by answer-key instability and representation-level specification errors, with pronounced disciplinary heterogeneity in defect structure.
@@
    Incorrect final answers dominate answer-level instability. Deterministic answer-key errors constitute the principal failure mode across all subjects.
@@
    Rationale instability is primarily structural or representational. Missing intermediate steps and format-level semantic misalignment outweigh classical logical contradictions.
@@
    Representation sensitivity is discipline-dependent. Format semantic errors play a particularly large role in Chemistry and Computer Science.
@@
    Conceptual invalidity is not the dominant driver. Most defects arise from specification or answer-key instability rather than from fundamentally incorrect task concepts.
These findings suggest that defect patterns in HLE exhibit systematic structure rather than uniform random noise. Reliability risks are component-specific and discipline-sensitive, with answer correctness and rationale completeness constituting the dominant bottlenecks. This empirical structure justifies a component-wise verification protocol and cautions against treating benchmark noise as uniform or negligible.
@@
Physics Semantic Error (57.1%) Information Missing (25.0%) Incorrect Answer (80.0%)
Chemistry Format Semantic Error (56.5%) Format Semantic Error (40.0%) Incorrect Answer (71.4%)
Biology/Medicine Knowledge Error (33.3%) Information Missing (45.0%) Incorrect Answer (97.2%)
Computer Science Format Semantic Error (94.6%) Format Semantic Error (63.3%) Incorrect Answer (78.1%)
@@ 5.1 Experimental Setup
We evaluate the following frontier models: GPT-5.2-Thinking (OpenAI, 2025), Gemini3-Pro-Preview (Google DeepMind, 2025), Claude-Opus4.5 (Anthropic, 2025), Claude-Opus4.6 (Anthropic, 2026), Grok-4.1 (fast-reasoning) (xAI, 2025), DeepSeek-V3.2-Thinking (Liu et al., 2025), Qwen3-Max-Thinking (Qwen, 2026a), Qwen3.5-Plus (Qwen, 2026b). All models were evaluated using the same system prompt recommended by the HLE official guidelines, along with each model’s own default recommended decoding configuration. To reduce variance from stochastic decoding, we run five independent rollouts per item and report avg5 accuracy, i.e., the average correctness across the five sampled completions (without additional post-hoc selection unless otherwise stated).
@@
We report results on (i) the Full Set (Raw HLE Full vs. Revised HLE-Verified Full) to quantify the end-to-end impact on the benchmark, and (ii) a Revised Subset comparison (Raw HLE Subset vs. Revised HLE-Verified Subset) limited to items that were edited/flagged during our verification process, which isolates the effect of flawed items and their corrections. Note that, to ensure comprehensive revisions, we also corrected the original rationales and released them to the open-source community. However, under the standard HLE benchmark evaluation, items with errors only in the rationale do not affect the final scores, since mainstream evaluation uses only the problem and answer. Therefore, the Revised Subset in this experimental section includes only items for which at least one of the problem/answer fields was revised. Unless otherwise specified, all evaluation sets used in this section are the text-only subset of HLE.
@@
Claude-Opus4.5 30.00 55 38.8 46 8.80 14.20 70 48.16 39 33.96
Grok 4.1-fast-reasoning 19.94 73 29.0 63 9.06 8.25 83 43.07 47 34.82
@@
Deepseek-V3.2 24.90 56 36.4 46 11.50 7.87 70 47.45 28 39.58
Qwen3-Max-Thinking 30.00 66 38.2 57 8.20 14.2 79 48.48 44 34.28
@@ A.1 LLM Judge Prompt
A.1.1 LLM Judge Prompt in Stage I
@@
A.1.2 LLM Judge Prompt in Stage II